OpenStudy (anonymous):

Some please help Harry can rake the leaves in the yard 8 hours faster than his little brother Jimmy can. If they work together, they can complete the job in 3 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jimmy to complete the job on his own.

OpenStudy (dumbcow):

use rates, the rate a person is working is the reciprocal of the time working together, they work at a rate of 1/3 job per hour when two people work together their rates add together ---> 1/h + 1/j = 1/3 ---> h = j - 8 (harry takes 8 hrs less than jimmy) substitute and solve for j \[\frac{1}{j-8} + \frac{1}{j} = \frac{1}{3}\] combine fractions \[\frac{j+(j-8)}{j(j-8)} = \frac{1}{3}\] cross-multiply \[j^2 -8j = 6j -24\] \[j^2 -14j +24 = 0\] Factor \[(j-12)(j-2) = 0\] j = 12 It takes jimmy 12 hours to rake the leaves